OpenStudy (anonymous):

(5{sqrt}6-10{sqrt}5)(3{sqrt}6+6{sqrt}5) multiply

OpenStudy (anonymous):

\[(5\sqrt{6}-10\sqrt{5})(3\sqrt{6}+6\sqrt{5})\]